OpenStudy (anonymous):

Let u = <5,6>, v = <-7,2>. Find 8u - 5v.

OpenStudy (unklerhaukus):

\[X=\langle x_1,x_2\rangle \] \[\alpha X=\alpha\langle x_1,x_2\rangle=\langle \alpha x_1,\alpha x_2\rangle\]

OpenStudy (unklerhaukus):

\[Y=\langle y_1,y_2\rangle\]\[Z=\langle z_1,z_2\rangle\] \[Y+Z=\langle y_1,y_2\rangle+\langle z_1,z_2\rangle=\langle y_1+z_1,y_2+z_2\rangle\]